OpenStudy (anonymous):

Simplify sqrt -72 ??

OpenStudy (anonymous):

\[\sqrt{-72}\] ANSWER CHOICES: \[6\sqrt{-2}\] \[-6\sqrt{2}\] \[6\sqrt{2i}\] \[6i \sqrt{2}\]

OpenStudy (anonymous):

ALSO, Simply i^41 ANSWER CHOICES: -1 -i 1 i

OpenStudy (lgbasallote):

woww..two questions in one post...

OpenStudy (anonymous):

@lgbasallote lol yes i know

OpenStudy (anonymous):

the last choice

OpenStudy (lgbasallote):

\[\sqrt{-72} \implies \sqrt{72 \times -1} \implies \sqrt{72} \times \sqrt{-1}\] if \[\sqrt{72} \implies 6\sqrt 2\] and \[\sqrt{-1} = i\] the what would be \[\sqrt{72} \times \sqrt{-1}\] does that help?